OpenStudy (anonymous):

Find the y value for the point that splits segment GH in half if point G is located at (−3, 5) and point H is located at (0, −2). 4 2 3.5 1.5

OpenStudy (anonymous):

Is there a graph to go along with this?

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

take the average

OpenStudy (anonymous):

Ok, I think I can solve this or you can go with satellite73

OpenStudy (anonymous):

\[\frac{5-2}{2}=\frac{3}{2}=1.5\]

OpenStudy (anonymous):

since you just want the \(y\) value the point itself is \[(\frac{-3}{2},\frac{3}{2})\]

OpenStudy (anonymous):

Satellite73 has it, you would pick the last choice

OpenStudy (anonymous):

Thank you both so much!